. FV of $400 paid each 6 months for 5 years at a nominal rate of 12%

    compounded semiannually


1
Expert's answer
2021-06-15T13:45:50-0400

"FV = PMT \frac{(1+ \frac{r}{n})^{t \times n}-1}{\frac{r}{n}}"

PMT=$400

r=12%

t=5

n = 2

"FV = 400 \times \frac{(1+ \frac{0.12}{2})^{5 \times 2}-1}{\frac{0.12}{2}} \\

= 400 \times \frac{1.7908477 -1}{0.06} \\

= 5272.32"